# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ V-C Intent-based Networking
Intent-based networking (IBN) is a network management paradigm that translates high-level business objectives, or human-readable “intents”, into automated network configurations [176]. Through closed-loop automation, IBN continuously verifies and enforces these intents, focusing on three core processes [177]: (i) Intent refinement, which involves the translation of user-defined intents into maintainable and adaptable network policies, (ii) Intent activation, which uses network analytics and machine learning to identify and resolve policy conflicts, and (iii) Intent assurance, which integrates with SDN controllers and orchestration systems, using telemetry to continuously monitor network functions and ensure that intents are met through ongoing adjustments.
@@
The effectiveness of IBN depends heavily on the accurate and consistent interpretation of user inputs, such as voice, text, or other data, and the extraction of intents from these inputs. As illustrated in Fig. 12, TinyML–LargeML integration supports this goal. TinyML serves as a compact node for real-time, on-device AI processing, enabling accurate reflection of user intents and facilitating coordinated network behavior [178]. LargeML functions as a core server, aggregating intents from TinyML nodes using techniques such as TL, FTL, FFMs, SL, FSL, or FSFMs, and performing subsequent translation into actions and policies without human intervention [179]. A notable example of TinyML–LargeML synergy in IBN is an intent-based on-device AI framework [180] that facilitates collaborative AI across resource-constrained IoT devices by capturing and reflecting user intents. In this framework, TinyML manages local processing of multi-modal sensor data, for example speech-to-text conversion and an intent translation, while LargeML centrally manages policy coordination. Reducing dependence on cloud-based processing improves security, processing speed, and energy efficiency. TinyML–LargeML integration for IBN has also been applied to resource-constrained autonomous aerial vehicle swarm networks [181]. This framework combines TinyML with explainable AI to support the tamper-proof security features of blockchain and zero-trust architecture, enabling time-sensitive intent-based predictions with interpretable cognitive intelligence. It allows the detection of false communications that could distort positional data shared between aerial vehicles. The numerical results in [181] indicate a 2.34% increase in prediction accuracy, a 2.97% gain in sensitivity, latency reduction to 43.1 ms, and storage reduction to 0.0013 MB.
@@ V-D Zero-touch Network
The zero-touch network (ZTN) is a management paradigm that uses AI, ML, and IBN to automate operations with minimal to no human input [182]. ZTN improves network efficiency by automating routine tasks such as configuration, monitoring, and troubleshooting, which lowers operational costs and frees administrators to focus on strategic planning. ZTN also improves transmission reliability by continuously analyzing traffic and user requests, adjusting network parameters in real time, resolving faults, and restoring services autonomously. ZTN also aligns network behavior with organizational policies while maintaining the flexibility to scale with increasing network demands and evolving requirements. Importantly, ZTN supports real-time threat detection and protection by integrating advanced security mechanisms.
@@
Achieving zero-touch automation with zero net emissions in AI-driven wireless networks requires energy-efficient TinyML–LargeML integration, low-power transceivers, lightweight AI models, and adaptive resource allocation [183]. In response, research on energy-efficient AI has accelerated, and Fig. 13 is a vision of the integrated TinyML–LargeML ecosystem for 6G ZTNs. To support autonomous solutions in self-organizing networks, TinyML optimizes neural network training on IoT devices for sustainable local learning, and LargeML uses FL and DL at base stations or clouds to reduce energy use and balance computational load [184]. In mobile networking, TinyML enables autonomous IoT node operations, and LargeML supports coverage map generation and edge AI-as-a-service (AIaaS), as seen in NanoDeploy Automator [185]. In industrial IoT, TinyML extracts data at terminal nodes, with LargeML at gateways and clouds enhancing efficiency and traffic management, exemplified by Open RAN (O-RAN)-based ZTNs for smart factories [186, 187]. Cross-layer cybersecurity frameworks like SH-CASH use AutoML to secure physical and protocol layers, with potential for TinyML integration within O-RAN architectures, ensuring service-level agreement (SLA) compliance and high performance [188].
@@ V-E Brain-level Metaverse
The Brain-Level Metaverse (BrainMeta) integrates brain-computer interface (BCI) technology with the metaverse to create immersive virtual environments where users control avatars, manipulate objects, and navigate using brain signals [189]. Enabled by non-invasive (e.g., EEG headsets) [190] or invasive (e.g., brain implants) [191] BCI methods, BrainMeta adapts to users’ cognitive states, delivering personalized experiences for applications like virtual learning, gaming, remote work, and rehabilitation therapies for neurological conditions. It also opens employment opportunities for individuals with disabilities by supporting metaverse-based workspaces.
@@
Despite its potential, BrainMeta remains in an early developmental stage and faces key challenges, particularly in building brain-inspired cognitive architectures constrained by size, weight, power, and processing limits [192]. BrainMeta also demands more efficient, high-performance models to accurately interpret and represent information for enhanced user interaction, cognition, content creation, and service quality in alignment with neural characteristics. These challenges can be negotiated through the integrated deployment of TinyML and LargeML, as illustrated in Fig. 14. TinyML algorithms can exploit neuromorphic computing chips [193] to support diverse message formats across computing processes. This capability enables customization of processing elements in brain-inspired communication devices, localizes memory, and increases communication bandwidth, enhancements which accelerate image recognition and classification tasks. In addition, TinyML can operate as an edge-based platform within the metaverse, using efficient TinyML processors to reduce latency and resource demands [194]. One such optimization involves winner-take-all circuits implemented with simplified leaky-integrate-and-fire neurons on field-programmable gate arrays. LargeML-based approaches [9], such as cloud–edge–end collaborative models, mobility-aware pre-rendering, and diffusion model-based adaptive rendering, can be improve data collection from TinyML devices and enhance rendering tasks in the metaverse. The process involves extraction of generalized representations from extensive EEG data, segmentation of data into channel arrays, and application of vector-quantized neural spectrum prediction to train and update neural analyzers [195]. AMFL, an adaptive, resource-efficient FL framework designed for metaverse environments, offers another example [196]. By combining FL and DRL strategies within BrainMeta, AMFL mitigates the adverse effects of non-IID data, reduces resource demands, and improves the QoE for human-centric, resource-constrained augmented reality devices.
@@ VI-A Standardization
5G faces challenges with dynamic user demands and traffic, while 6G aims to enable sustainable applications like holography and the metaverse through TinyML–LargeML integration [197]. This integration is designed for user-centric services that adapt to changing behaviors and networks. However, standardization issues impede the development of a fully integrated ecosystem. The ITU-R identifies RAN slicing as crucial for 6G virtual network sharing [198], requiring a unified O-RAN architecture that supports diverse air interfaces and AI functions. This architecture should offer customizable service parameters, but shared infrastructure raises concerns about transparency, reliability, and privacy. Standardization bodies are advancing AI-driven wireless architectures. 3GPP’s Releases 17 [199] and 18 [200] enhance 5G RAN with AI for intelligence and efficiency, while the O-RAN Alliance integrates AI for orchestration and control [201, 202].
@@ VI-B Resource Management and Orchestration
6G networks face resource management and orchestration challenges, including wireless channel uncertainty, ISAC device proliferation, and carbon emission reduction. These drive AI/ML-based frameworks to optimize spectrum, power, and time allocation, enhancing interference management, energy efficiency, and reliability over traditional methods like channel estimation and beamforming [207]. However, integrating AI/ML introduces data privacy risks, computational overhead, and scalability issues. Adaptive algorithms like DRL with diffusion models [207] and distributed learning like FTL with blockchain incentives [208] can mitigate these by preserving privacy and reducing latency.
Deploying DL with TinyML faces obstacles [209]: (i) computational trade-offs, (ii) vanishing gradients needing hardware acceleration, (iii) memory constraints for semantic communication, (iv) reliance on scarce real-world data, and (v) dynamic environmental variability. LargeML, using distributed neural networks and collaborative learning, encounters issues like: (i) data segmentation for collective learning, (ii) transmission errors in FL affecting accuracy, and (iii) stringent privacy and coordination demands. Future research into model-agnostic meta-learning [210] or accretionary learning [211] could address these challenges.
At a broader level of cross-domain interaction and orchestration, AI will be integrated into 6G network slicing, leveraging NFV for resource virtualization and SDN for centralized management. A central SDN controller handles planning and provisioning, while AI-enhanced local controllers manage operations, ensuring cost-effective, QoS-compliant network slices. However, cross-domain AI coordination poses challenges, including those outlined in [212] as follows: (i) Data/knowledge coordination, (ii) model training coordination, (iii) inference and decision coordination, and (v) computing resource coordination.
Beyond technical integration, managing and orchestrating TinyML and LargeML in 6G network slices is challenging due to the shift from 5G’s centralized learning to distributed paradigms like FL, swarm learning, multi-intelligent RL, TL, meta-learning, and SL. The choice of learning approach depends on node capabilities, dataset traits, computing resources, model complexity, latency, and network conditions. Supporting diverse distributed methods increases orchestration complexity. To address this, AI nodes should adopt specialized roles, such as horizontal FL for decentralized data, hierarchical FL for reduced communication, TL for efficient training, vertical FL for cross-domain learning, SL for uneven resources, multi-intelligent RL for localized collaboration, and swarm learning for secure joint training.
@@ VI-C Advanced Security and Privacy-Preserving Techniques
As TinyML–LargeML integration becomes central to next-generation wireless network developments, the security and privacy landscape will continue to grow in complexity. Key concerns include adversarial attacks, data security, model trustworthiness and integrity, and user privacy.
TinyML faces security challenges due to limited edge device resources, making devices like traffic sensors, smartphones in FL, and wearables vulnerable to adversarial attacks. Adversarial training [213] enhances TinyML’s resilience to malicious inputs, while homomorphic encryption [214] ensures data confidentiality and integrity. Privacy risks from sensitive edge data can be mitigated using differential privacy, which adds noise to obscure information [215], and advanced encryption to secure model parameters during transmission [216].
LargeML deployment also requires robust privacy and security due to its broader scope and multi-stakeholder involvement. Model distillation [33] transfers LargeML knowledge to TinyML, reducing computational needs while maintaining performance. Integrating LargeML with GANs [217] simulates cyberattacks to improve resilience. Blockchain ensures model integrity by hashing and recording versions, maintaining TinyML model authenticity [215]. Advanced privacy-preserving techniques further secure LargeML systems: (i) privacy-preserving data shuffling, (ii) differential private synthetic data generation, iii) private aggregation of teacher ensembles, (iv) Rényi differential privacy for generative modeling, (v) cross-Silo FL with secure data sharding, and (vi) decentralized privacy-preserving knowledge distillation.
@@ VI-D Real-time, Lightweight Intelligence
Beyond 5G and 6G networks aim to automate M2M operations by transforming conventional communication systems into AI-centric architectures, ultimately transitioning toward human-centric networks through advanced AI integration. To meet the growing demands for improved user experience and reduced carbon emissions, the integration of TinyML and LargeML into AI-enabled systems creates new opportunities for real-time, lightweight intelligence yet also introduces distinct technical challenges.
TinyML applications vary in communication protocols, latency requirements [11], and device specifications. For example, Arduino Nano (64 MHz, 1 MB flash memory, 256 KB RAM) and Raspberry Pi (up to 1.8 GHz, 32 GB flash memory, 8 GB RAM) reflect a range of processing capabilities [4]. This variability highlights the need for compact, unified TinyML frameworks that support flexible plug-and-play deployment, provide efficient processing, and maintain lightweight runtimes [218]. In contrast to TinyML, LargeML typically processes vast input volumes from multiple domains, each requiring tailored prompt engineering and model-splitting approaches. To deliver real-time responses with accurate and reliable outputs, LargeML often depends on extensive hyperparameter tuning and multi-step processing. These operations result in prolonged training times, high power consumption, and significant memory requirements [16]. Consequently, future LargeML should adopt adaptive prompt engineering strategies, such as model-adaptive prompts [219] and lightweight tuning paradigms for pluggable prompting [220], to improve explainability and accelerate processing and inference without disruption to the life cycle pipeline.
However, the limited availability of unified platforms and tools for managing workflows, coordinating input data, and ensuring target environment compatibility poses additional challenges in TinyML–LargeML integration. For example, LLMs require specialized ML libraries, compilers, and runtimes for code generation, text processing, and prompt execution tasks, each with varying input formats, volumes, and execution times [25]. Iterative error-handling and refinement mechanisms are often required to improve output reliability. Thus, in addition to concentrating on optimizing TinyML and LargeML design, comprehensive guidelines and evaluation frameworks are essential to define task-specific roles and system requirements for each deployment and operation phase [221].
VI-E AI-Native and Collective AI for 6G
The concept of AI-Native redefines system design by positioning AI at the core of operations rather than treating it as supplementary feature, as seen in autonomous vehicles where AI manages navigation, obstacle detection, and collision avoidance. As defined in [222], future AI-native systems aim to unify five critical capabilities: solving complex problems via data-driven AI for seamless operations, adapting to dynamic environments, combining foundational knowledge with new insights for continuous learning, ensuring trustworthy and transparent AI functions, and enabling resilient, cognitive autonomous networks. Nevertheless, integrating TinyML and LargeML into AI-Native systems poses challenges: (i) coordinating resource-constrained TinyML with computationally robust LargeML for zero-touch operations, (ii) defining their roles across heterogeneous infrastructures to have a unified framework, (iii) ensuring interoperable learning to support continuous learning across the network, (iv) managing AI life cycles for reliability and fairness, and (v) optimizing joint performance while maintaining TinyML accuracy.
Collective AI involves multi-agent collaboration, where agents contribute to shared goals through local training without direct communication [223]. At the platform level, collective reinforcement learning enables agents to share models, thereby achieving short training time, redundant learning avoidance, and robust models from limited samples [224]. At a broader implementation level, collective AI swarm intelligence supports IoT applications in static and dynamic environments [225], but it requires designing a unified framework for collective dynamics and reasoning across heterogeneous systems.
As the integration of TinyML and LargeML progresses within collective AI systems, new opportunities will emerge alongside specific challenges: (i) Crowd retention: Varying agent contributions require robust incentive mechanisms to sustain participation; (ii) Spectrum utilization: Dense IoT networks strain spectrum resources, necessitating optimized bandwidth allocation, especially with 6G; (iii) Security, trust, and privacy: Sharing training outputs raises concerns, with blockchain’s resource demands limiting its use on constrained devices, while human-AI collaboration needs better social cue interpretation, requiring advanced cognitive architectures; and (iv) Ethical issues: Balancing TinyML and LargeML benefits with risks involves addressing transparency, bias, and regulatory gaps. Errors in LargeML outputs could have significant consequences, underscoring the need for legal frameworks to ensure accountability and trust.
VII Conclusion

### v2

@@ V-C Intent-based Networking
Intent-based networking (IBN) is a management paradigm that translates high-level business objectives, or human-readable “intents”, into automated network configurations [32]. Using closed-loop automation, IBN continuously verifies and enforces these intents through three processes [124]: (i) Intent refinement, where user-defined intents are translated into maintainable and adaptable network policies; (ii) Intent activation, which leverages analytics and ML to detect and resolve policy conflicts; and (iii) Intent assurance, which integrates with SDN controllers and orchestration systems, using telemetry for continuous monitoring and adjustment.
@@
The effectiveness of IBN depends on the accurate interpretation of user inputs (e.g., voice, text, multi-modal data). As shown in Fig. 12, TinyML–LargeML integration supports this objective. TinyML provides lightweight, on-device AI for real-time intent detection and localized processing [174], while LargeML aggregates intents across devices and translates them into network actions and policies using techniques such as TL, FTL, FFMs, SL, FSL, and FSFMs [112]. A notable example is an intent-based on-device AI framework [3], where TinyML processes local sensor data (e.g., speech-to-text conversion and intent translation), while LargeML manages centralized policy coordination. This reduces reliance on cloud-only processing, improving security, latency, and energy efficiency. Another application is intent-based management for autonomous aerial vehicle swarm networks [6], which combines TinyML with explainable AI, blockchain, and zero-trust architectures. This system supports interpretable, tamper-proof intent predictions and detects false communications that may distort shared positional data. Results in [6] demonstrate a 2.34% accuracy gain, 2.97% sensitivity improvement, latency reduced to 43.1 ms, and storage savings down to 0.0013 MB.
@@ V-D Zero-touch Network
A zero-touch network (ZTN) leverages AI, ML, and IBN to automate operations with minimal human intervention [109]. By handling tasks such as configuration, monitoring, and troubleshooting, ZTNs reduce operational costs, improve efficiency, and free administrators for strategic planning. They also enhance reliability by analyzing traffic and user requests in real time, adjusting parameters, resolving faults, and restoring services autonomously. Moreover, ZTNs align network behavior with organizational policies while adapting to growing demands and evolving requirements. Importantly, ZTNs integrated security mechanisms further enable real-time threat detection and protection.
@@
Achieving zero-touch automation with net-zero emissions in AI-driven wireless networks demands energy-efficient TinyML and LargeML integration, low-power transceivers, lightweight AI models, and adaptive resource allocation [2]. Recent research highlights a TinyML-LargeML ecosystem for 6G ZTNs (see Fig. 13). In self-organizing networks, TinyML optimizes neural network training on IoT devices, while LargeML uses federated and deep learning at base stations or clouds to balance computational loads and minimize energy consumption [152]. In mobile networking, TinyML allows autonomous IoT operations, while LargeML develops coverage maps and offers AI-as-a-service (AIaaS), as shown in NanoDeploy Automator [147]. In industrial IoT, TinyML processes data locally, while LargeML enhances efficiency and traffic management at gateways and clouds, as seen in O-RAN-based ZTNs for smart factories [182]. Cross-layer cybersecurity frameworks, like SH-CASH, utilize AutoML to secure layers, with TinyML ensuring service-level agreement compliance in O-RAN architectures [196].
@@ V-E Brain-level Metaverse
The brain‑level metaverse (BrainMeta) links brain computer interfaces (BCIs) with immersive virtual worlds, allowing users to control avatars, interact with objects, and navigate environments through brain signals [211]. BCIs range from non‑invasive EEG headsets [75] to invasive implants [36], enabling personalized experiences in virtual learning, gaming, remote work, and neurorehabilitation while expanding accessibility for users with disabilities. Yet BrainMeta remains early‑stage and must overcome challenges in designing brain‑inspired cognitive architectures under strict size, weight, power, and processing limits [201], as well as the need for efficient, high‑performance models for neural signal interpretation, cognition support, and adaptive content generation, motivating integrated TinyML–LargeML deployment (see Fig. 14).
TinyML can leverage neuromorphic chips [82] to support diverse neural message formats, customize processing in brain‑inspired devices, localize memory, and accelerate tasks such as image recognition and classification, while edge‑based TinyML processors reduce latency and resource use through circuits like winner‑take‑all implementations on FPGA hardware [76]. LargeML complements this with cloud‑edge‑end collaborative models, mobility‑aware pre‑rendering, and diffusion‑based adaptive rendering [183], enhancing data collection from TinyML devices and improving metaverse quality. Techniques include extracting generalized representations from large‑scale EEG datasets, segmenting signals into channel arrays, and applying vector‑quantized neural spectrum prediction for training neural analyzers [74]. Additionally, AMFL [133] combines FL and DRL to handle non‑IID data, reduce resource overhead, and improve QoE for resource‑constrained, human‑centric AR devices.
@@ VI-A Standardization
5G faces challenges with dynamic user demands and traffic, while 6G aims to enable sustainable applications like holography and the metaverse through TinyML–LargeML integration [198]. This integration is designed for user-centric services that adapt to changing behaviors and networks. However, standardization issues impede the development of a fully integrated ecosystem. The ITU-R identifies RAN slicing as crucial for 6G virtual network sharing [188], requiring a unified O-RAN architecture that supports diverse air interfaces and AI functions. This architecture should offer customizable service parameters, but shared infrastructure raises concerns about transparency, reliability, and privacy. Standardization bodies are advancing AI-driven wireless architectures. 3GPP’s Releases 17 [164] and 18 [165] enhance 5G RAN with AI for intelligence and efficiency, while the O-RAN Alliance integrates AI for orchestration and control [99].
@@ VI-B Resource Management and Orchestration
6G networks face rising orchestration and resource management pressures due to wireless channel uncertainty, dense ISAC deployments, and carbon‑reduction requirements. AI/ML improves interference control, energy efficiency, and reliability [40], but also raises concerns about privacy, computational load, and scalability. New directions include DRL paired with diffusion models for adaptive optimization [40] and FTL with blockchain‑based incentives for low‑latency, privacy‑preserving coordination [184]. TinyML remains constrained by compute trade‑offs, vanishing gradients that demand hardware acceleration, memory limits for semantic communication, limited real‑world data, and environmental variability [163]. LargeML, despite distributed learning, still struggles with data segmentation, transmission errors in FL, and strict privacy and coordination needs. Future methods such as model‑agnostic meta‑learning [49] and accretionary learning [187] may help alleviate these challenges.
At the orchestration layer, integrating AI into 6G network slicing blends NFV‑based virtualization with SDN’s centralized control: a global SDN controller handles planning and provisioning, while AI‑enhanced local controllers optimize slice‑level performance. Cross‑domain AI coordination introduces challenges in data and knowledge sharing, training alignment, inference synchronization, and resource allocation [92]. As 6G shifts from centralized 5G models to distributed learning, learning-model orchestration becomes more complex and depends on node capacity, data properties, resource availability, model size, latency, and network dynamics. Assigning specialized roles across slices enables scalable, efficient learning in 6G environments, e.g., horizontal FL for decentralized data, hierarchical FL for reduced communication, TL for efficient reuse, vertical FL for cross‑domain learning, SL for heterogeneous resources, multi-agent RL for local coordination, and swarm learning for secure joint training.
@@ VI-C Advanced Security and Privacy-Preserving Techniques
As TinyML–LargeML integration becomes central to next‑generation wireless networks, security and privacy concerns intensify, especially around adversarial threats, data protection, model integrity, and user privacy. TinyML’s limited edge‑device resources make sensors, smartphones in FL, and wearables vulnerable to attacks, but adversarial training strengthens robustness [83], homomorphic encryption protects data confidentiality and integrity [172], differential privacy obscures sensitive information [8], and advanced encryption secures model parameters during transmission [5]. LargeML, operating across broader and multi‑stakeholder environments, also requires strong safeguards: model distillation reduces computational demands while preserving performance [194], GAN‑based simulations improve cyber‑resilience [177], and blockchain ensures model integrity by hashing and tracking TinyML model versions [8]. Additional privacy‑preserving techniques, including data shuffling, machine unlearning [122], synthetic data generation, private teacher‑ensemble aggregation, Rényi differential privacy, secure cross‑silo FL, and decentralized knowledge distillation, further reinforce LargeML deployments.
@@ VI-D Real-time, Lightweight Intelligence
Beyond‑5G and 6G networks are shifting from traditional M2M automation toward human‑centric architectures, demanding AI systems that provide real‑time, lightweight intelligence while meeting user‑experience and sustainability goals. Joint TinyML–LargeML deployment expands these capabilities but also introduces challenges. TinyML devices span a wide range of protocols, latency profiles, and hardware, from Arduino Nano (64 MHz, 1 MB flash, 256 KB RAM) to Raspberry Pi (1.8 GHz, 32 GB flash, 8 GB RAM) [1, 97], emphasizing the need for compact, unified frameworks that support plug‑and‑play deployment, efficient processing, and lightweight runtimes [65]. LargeML, by contrast, handles massive multimodal inputs and depends on prompt engineering, model partitioning, and hyperparameter tuning, all of which incur high training costs, power consumption, and memory demands [135]. Emerging techniques such as model‑adaptive prompts [27] and pluggable lightweight tuning [25] can improve explainability and accelerate inference without disrupting life‑cycle pipelines.
Integration remains limited by the lack of standardized platforms and tools for workflow management, data coordination, and runtime compatibility. Large models often require specialized libraries, compilers, and runtimes with inconsistent input/output specifications [189], making iterative refinement necessary for reliable outputs. As a result, progress depends not only on optimizing TinyML and LargeML individually but also on developing guidelines and evaluation frameworks that define task‑specific roles, interoperability requirements, and life‑cycle management strategies across domains [59].
VI-E AI‑enabled 6G Communication-Sensing-Computing
Integrating TinyML’s real‑time, low‑power intelligence with LargeML’s high‑capacity reasoning creates a unified AI layer for 6G, where communication, sensing, and computation operate as a single adaptive system [110]. In this hierarchy, TinyML provides fast on‑device perception through lightweight feature extraction and inference, while LargeML offers deeper interpretation and coordination through global optimization, multimodal fusion, and long‑horizon prediction. To do so, TinyML first semantically compresses sensed data for efficient transmission [203], LargeML then uses these representations to optimize routing, beamforming, and spectrum allocation, and finally, LargeML’s predictions guide TinyML nodes to adjust sensing rates or modalities, forming adaptive networks responsive to mobility, interference, and environmental changes. This synergy supports scalable multimodal sensing, energy‑efficient operation, and human‑centric services [46], where edge devices capture fine‑grained context and cloud models infer intent or anticipate user needs.
However, integrating TinyML and LargeML in 6G introduces significant challenges. First, model partitioning must balance latency, energy, and accuracy across heterogeneous devices [203]. Second, synchronizing TinyML’s millisecond‑scale inference with LargeML’s slower global reasoning requires new timing and scheduling mechanisms. Third, interoperability is hindered by diverse hardware, runtime environments, and data formats across edge devices [108]. Fourth, security and privacy risks increase as sensitive edge data interacts with large‑scale models, demanding robust protection against adversarial attacks, leakage, and model manipulation. Finally, additional 6G‑specific constraints (i.e., URLLC requirements, dynamic spectrum conditions, mobility‑induced channel variation, and sustainability targets [110, 46]) further complicate joint deployment.
VI-F AI-Native and Collective AI for 6G
AI‑native systems embed intelligence at the core of their operation, aiming to solve complex tasks through data‑driven reasoning, adapt to dynamic environments, fuse foundational knowledge with new insights, and maintain transparency and trustworthiness for resilient autonomous networks [69]. Integrating TinyML and LargeML into these architectures introduces challenges such as coordinating constrained TinyML with powerful LargeML for zero‑touch automation, defining their roles across heterogeneous infrastructures, ensuring interoperable lifelong learning, managing AI life cycles for fairness and reliability, and balancing joint performance without degrading TinyML accuracy. Collective AI extends this vision through multi‑agent collaboration, where agents learn locally toward shared goals without direct communication [209]; platform‑level collective reinforcement learning accelerates convergence and reduces redundant training [90], while system‑level swarm intelligence supports IoT applications in dynamic environments [145], though it still lacks unified reasoning and coordination frameworks.
Integrating TinyML and LargeML into collective AI creates new opportunities but also raises several challenges: sustaining agent participation through incentive mechanisms, ensuring efficient spectrum use in dense 6G IoT deployments, and addressing security, trust, and privacy concerns as shared outputs may expose sensitive information while blockchain remains resource‑intensive for constrained devices. Human-AI collaboration further requires cognitive architectures capable of interpreting social signals, and ethical considerations, including transparency, bias mitigation, and accountability, are essential, especially as errors in LargeML outputs can have significant real‑world consequences.
VI-G Application Prospects and Outlook for 6G and AI
6G will unify TinyML at the edge with large-scale AI models deployed across the network and cloud, creating an intelligent and adaptive wireless environment. This integration will enable real-time sensing, IBN, semantic communication, and automated SDN/NFV orchestration, thereby forming the foundation for immersive XR, holographic telepresence, and large-scale digital twins [67, 14]. Early scenarios already reflect this direction: in smart factories, TinyML sensors detect anomalies locally and send only semantic insights, while edge AI coordinates predictive maintenance and robotic actions with sub‑second latency. At the city scale, distributed TinyML nodes monitor environmental and public‑safety conditions and feed compact features into collective AI systems that coordinate drones, traffic flows, and emergency responses through privacy‑preserving FL.
Over the longer term, 6G is expected to evolve into an AI‑native network where intelligence is embedded across all layers, and devices collaborate through distributed learning and collective decision making [14]. This will enable ZTN operation, self‑healing infrastructures, and brain‑level Metaverse experiences that tightly integrate communication, sensing, and computation. Realizing this vision requires advances in O‑RAN standardization, trustworthy and explainable AI, secure FL, and efficient orchestration of communication, computing, and energy resources. As these foundations mature, the fusion of 6G and AI will support resilient industrial automation, new service models, and globally scalable intelligence.
VII Concluding Remarks